# Time Series Information Visualization – A Review of Approaches and Tools

### v1

@@ 8 Conclusion
Visualization has the power to transform abstract phenomena into understandable visual elements [160]. In this review, we have explored the role of visualization systems in making the implicit information hidden in time series data both explicit and comprehensive. However, several factors can influence the effectiveness of visual tools in communicating information, including the selection of appropriate graphical metaphors, the implementation of interactive data manipulation techniques, user-centric designs that retain human cognitive abilities, and even the quality and dimensions of devices used to access information [169]. Addressing these factors is crucial for maximizing the potential of visualization in data interpretation.
Time dependency can make the analysis of complex data even more challenging. A well-designed presentation of time series data is fundamental for accurate interpretation, as its structure is inherently complex and often difficult to manipulate in the raw state. Moreover, conventional data formats (such as large spreadsheets) are not straightforward to handle on a large scale and can obscure data scientists’ discovery of relevant time-dependent patterns and trends. Therefore, it is fundamental to develop visual tools that provide clear and comprehensible layouts, allowing analysts to extract meaningful and valuable information from complex time series data.

### v2

@@ 8 Conclusion
The analysis of temporal dependencies in large-scale, multidimensional time series datasets presents significant challenges in many application domains, as evidenced by the references discussed in this review. Visualization has the power to transform abstract phenomena into understandable visual elements [171]. Therefore, well-designed visualization is fundamental in transforming abstract, time-oriented phenomena into perceptually intuitive representations, enabling analysts to uncover hidden insights with clarity. In Sections 5 and 6, we explored the role of visual approaches and methods that address the complexity of time series data to make implicit information explicit and comprehensive.
Several factors can impact the effectiveness of visualization tools in communicating information, including the selection of appropriate graphical metaphors (see Section 5), the implementation of interactive manipulation techniques (see Section 4.8), user-centric designs that align with cognitive capacities (see Section 6), and even the influence of device quality and dimensions on rendering fidelity [181]. Addressing these factors is essential for enhancing the potential of information visualization in data interpretation. Table 3 synthesizes visualization concepts exploited to mitigate the challenges related to visual data exploration, while Sections 7.2 and 7.3 outline open questions and future directions. Finally, we compared visualization software and environments to create InfoVis tools, highlighting their strengths for diverse applications (see Table 4). In particular, recent breakthroughs, such as generative visualization models, have achieved prominent results in improving the understanding, synthesis, and optimization of visual representations (see Section 6). Therefore, developing visualization tools that provide clear and comprehensible layouts in an era of ever-growing data complexity is fundamental to allow analysts to extract meaningful information.